OpenStudy (anonymous):

find the derivative of log3 e^x

OpenStudy (danjs):

can you rewrite that as ln(e^x) / ln(3)

OpenStudy (anonymous):

uhm idk..

OpenStudy (anonymous):

Look up the change of base formula for logarithms if you're not sure @dolphins121

OpenStudy (eric_d):

By using natural log..

OpenStudy (eric_d):

log(a)x=(lnx)/(lna)

OpenStudy (anonymous):

so ln(e^x)/ln(3) ?

OpenStudy (anonymous):

\[ \log_3(e^x)= x\log_3(e)\\ \]This is just \(cx\).

OpenStudy (jhannybean):

\[\log_a (x) = \frac{\log_b(x)}{\log_b(a)} ~~\sf [\text{this is the change of base formula}]\]\[\log_3(e^x) = \frac{\log_e(e^x)}{\log_e(3)}\]And remember, \(\log_e(e^x) = \ln(e^x)\)